If George adds his current age to the age he will be next year on this​ date, the sum is 35 years. How old will George be 11 years from​ today?
Answer by ankor@dixie-net.com(22740)

If George adds his current age to the age he will be next year on this​ date, the sum is 35 years.
:
find his present age
a = his present age
a + (a+1) = 35
2a = 35 - 1
2a = 34
a = 17 is his present age
:
How old will George be 11 years from​ today?
17 + 11 = 28 yrs